OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[\frac{ x^2 - 16 }{ x^2 - 8x + 16 }\]

OpenStudy (anonymous):

I have a feeling the 16s cancel out

OpenStudy (raden):

hint : x^2 - 16 = (x+4)(x-4) x^2 - 8x + 16 = (x-4)(x-4) now, which parts u can cancel it ?

OpenStudy (anonymous):

(x - 4) cancels out (x-4)?

OpenStudy (raden):

yes, so what your answer then

OpenStudy (anonymous):

it would be: x + 4 ----- x - 4 correct?

OpenStudy (raden):

correct, btw, do u understand how to factor out that ?

OpenStudy (anonymous):

I know how to factor out, but not algebraic ratios.

OpenStudy (raden):

ok, cool :)